Ticket Number: HEA-NOC/20050409-2 Ticket Status: OPEN
Ticket Type: unscheduled Resolver: HEA-NOC
Ticket Opened: 20050409 12:43 UTC+1 Problem Start: 20050409 12:13 UTC+1
Ticket Update: 20050409 12:43 UTC+1 Problem End:
Site/line: Digital Depot
The bluenet link from TCD to the Digital Depot has gone down. Cannot connec
to the Digital Depot Cisco 3550.
Client RIFTFS has lost connectivity to HEAnet.
20050409 12:26 EK Contacted the digital Depot NOC. No maintenance is
planned for the Depot this weekend. Digital Depot will see if local hands
can be arranged.
Contacted Dublin City Council Traffic department, all their fibre links
are working. Traffic has no record of any outages.
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20050409-2